# Beginner’s Charm: Beginner-Heavy Teams Are Associated With High Scientific Disruption

### v1

@@ Abstract
Teams now drive most scientific advances, yet the impact of absolute beginners—authors with no prior publications—remains understudied. Analyzing over 28 million articles published between 1971 and 2020 across disciplines and team sizes, we uncover a universal and previously undocumented pattern: teams with a higher fraction of beginners are systematically more disruptive and innovative. Their contributions are linked to distinct knowledge-integration behaviors, including drawing on broader and less canonical prior work and producing more atypical recombinations. Collaboration structure further shapes outcomes: disruption is high when beginners work with early-career colleagues or with co-authors who have disruptive track records. Although disruption and citations are negatively correlated overall, highly disruptive papers from beginner-heavy teams are highly cited. These findings reveal a “beginner’s charm” in science, highlighting the underrecognized yet powerful value of beginner fractions in teams and suggesting actionable strategies for fostering a thriving ecosystem of innovation in science and technology.
@@
Disruption $|$ Innovation $|$ Career age $|$ Beginner authors $|$ Team science
@@
Imagine a research team with a high fraction of ‘beginner’ authors—individuals with no prior publication. How likely is this team to disrupt its field? Intuitively, one might expect such teams to be unlikely engines of discovery. Scientific creativity rarely arises ex nihilo; theories of scientific and technological change emphasize discovery as an endogenous, cumulative process in which past knowledge enables future progress [1, 2, 3, 4, 5, 6, 7, 8]. In this process, extended training and accumulated expertise allow researchers to stand, in Newton’s phrase, “on the shoulders of giants” [9, 10, 11]. Indeed, combinatorial accounts of creativity emphasize such atypical recombinations of well-established knowledge components as the substrate of breakthroughs [12, 13]. Moreover, as fields advance, the “burden of knowledge” intensifies, requiring ever longer training to reach the frontier and leaving less time to extend it [14, 15]. By this logic, beginner authors—who have had little time to digest this burden—may be handicapped in their capacity for field-shifting innovation and disruption. If mastery is required before divergence, then beginner-heavy teams may lack the cognitive materials and experiential grounding to engineer novelty.
@@
Yet, there are equally compelling reasons to expect that beginner-heavy teams may, in fact, be well-positioned to disrupt scientific advancement. Youth is historically linked to novelty in science: they are less encumbered by entrenched paradigms, more open to new ideas, and more willing to take intellectual risks [16, 17]. Freed from institutionalized pressures—such as administrative burdens and funding constraints that often steer senior researchers toward conservative publication strategies [18, 19, 20, 21, 22, 23, 24, 25, 26]—beginners may pursue bold ideas more readily. Moreover, deep disciplinary expertise can harden into attachment, making it harder for experienced scientists to selectively “unlearn” prevailing assumptions and embrace radically new ones [27, 28, 29, 30]. From this perspective, beginners’ absence of enculturation may be a strength precisely because they have not yet fully assimilated the burden of knowledge. Indeed, many landmark discoveries originated as debut publications—such as Turing Awardee Yann LeCun’s early doctoral work, which laid the foundation for backpropagation [31, 32]. Extending this logic to the team level, beginner-heavy groups may be especially disruptive: their fresh cognitive repertoires, freedom from orthodoxy, and high risk tolerance may outweigh individual knowledge deficits, enabling breakthrough contributions more readily than in teams with fewer beginners.
@@
While recent studies have explored the aggregated effects of mean career age on team disruption [33, 34], the role of the very youngest contributors—beginner authors—remains overlooked. In particular, it remains unclear how the disruptivity of teams changes as the fraction of beginners increases, and what underlying team dynamics and knowledge integration mechanisms drive these outcomes. To address this gap, we conduct a large-scale study of more than 2828M articles published between $1971$ and $2020$ across multiple scientific domains. Our analysis draws on SciSciNet V2 [35, 36], one of the most comprehensive and widely used bibliometric data lakes. Using SciSciNet’s name-disambiguated author data [37], we calculate career age as the number of years since an author’s first publication [38, 39, 27], and classify authors into three career stages: beginner (0 years), early-career (1–10 years), and senior researchers (11 or more years). These thresholds reflect typical academic transitions, with the early stages encompassing graduate training and initial faculty appointments, and seniority being attained after a decade of publishing [40, 41].
@@
We find strong evidence for a universal and previously undocumented pattern of “beginner’s charm”: an increased presence of beginner authors in teams is associated with higher scientific disruption and innovation. This association holds across disciplines, time periods, and team sizes, suggesting it reflects a general feature of how science advances rather than a peculiarity of specific fields or historical moments.
@@ Results
Teams with higher beginner-author ratios are more disruptive and innovative
@@
As our primary team performance metric, we use the disruption score, originally developed to distinguish between consolidating and destabilizing contributions in science and technology [42]. This score ranges from 1-1 to +1+1, capturing papers that reinforce or shift scientific trajectories, respectively (see Materials and Methods). Since its introduction, the disruption score has been widely adopted and validated across domains [43, 35, 44, 45, 46, 47], providing a robust lens for answering our research queries.
@@
We find that papers with a higher proportion of beginner authors tend to be more disruptive: In the overall dataset, the beginner ratio is positively correlated with a paper’s disruption percentile (Pearson’s $r=0.11$, Holm-adjusted $P<10^{-6}$), while early-career ratio shows a small positive association ($r=0.0007$, $P<10^{-6}$), and senior-author ratio shows a negative association (r=0.07r=-0.07, $P<10^{-6}$).
@@
While the overall correlations are informative, they obscure the role of team size, since beginner ratios are mechanically constrained by the number of authors and disruption itself is known to vary with team size [43]. To address this, we stratify the analysis by team size and recompute the associations using binned author ratios (Figure 1). The results are strikingly monotonic across team sizes: beginner ratios show uniformly strong positive correlations with disruption (Kendall’s $\tau$ up to 1.0 with narrow 95% CIs, $P<0.001$ in larger teams), senior ratios show consistently negative associations (often $\tau=-1.0$, $P<0.001$), and early-career ratios display weaker but positive effects that reach significance in larger teams (Supplementary Table S1; all $P$-values are Benjamini–Hochberg FDR-corrected).
@@
Prior studies found that larger teams tend to produce less disruptive work [43], and we replicate this pattern in teams with low beginner ratios. As the beginner ratio increases, disruption levels rise across team sizes—and especially steeply so in large teams, where the usual size penalty attenuates (Supplementary Figure S1). In fact, beginner-heavy large teams approach the disruption levels of beginner-heavy small teams. One possible explanation is that a higher share of beginners may counteract the convergence pressures of large teams—such as reduced risk-taking [48], resistance to external perspectives [49], and neutralized viewpoints [50]—enabling a shift from consolidating “safe bets” to more exploratory, high-variance work [51, 52].
The positive association between beginner share and disruption is robust across all five decades and in all of 146 disciplines with at least 2000 data points in the corpus ($r>0$, $P<0.001$ throughout; Supplementary Tables S2-S3). Decade-specific analyses show the effect at every team size ($r>0$, $P<10^{-6}$ throughout; Supplementary Figure S2). Discipline-specific splits likewise show persistent positive associations for small teams (sizes $\leq 5$); for large teams, the trend is weaker in fields where beginner-heavy large teams are rare (Supplementary Figure S3). Collectively, these patterns indicate that disruption is robustly elevated in teams with a higher fraction of beginners.
@@
To verify that our findings are not an artifact of the disruption score metric, we employ a secondary metric of atypical combination scores (see Materials and Methods) [14, 53, 54, 55]. This score conceptualizes discovery as innovative knowledge-recombination: for each paper, all pairs of reference venues are scored for atypicality against a degree- and time-preserving null, and the median $z$-score of all these pairs is taken as the paper’s atypical combination score. A score below 0 represents innovative combinations of existing knowledge, and vice versa [53]. Unlike disruption (a downstream impact measure), the atypical combination score captures innovation from a paper’s inputs.
Figure 2 shows the cumulative distributions of atypical combination scores for different beginner author-ratio quartiles (Q1: 0–0.25, Q2: 0.25–0.50, Q3: 0.50–0.75, Q4: 0.75–1.00). As the share of beginner authors increases, the cumulative distributions shift toward lower values, indicating more innovative combinations. The ECDFs of the atypical combination scores differ significantly between every quartile pair (two-sample Kolmogorov–Smirnov test; Holm-adjusted $P<10^{-300}$ for all comparisons; largest contrast Q1 vs Q4: $D=0.121$). Supplementary Figure S4 shows that higher proportions of beginner authors are associated with lower atypicality score percentiles (higher innovation). In contrast, higher senior author ratios are associated with higher atypicality score percentiles (lower innovation). Together, these findings corroborate and validate our disruption-based results that beginner-heavy teams are more innovative than beginner-light ones.
The marked difference in disruption and innovation between beginner-heavy and other teams prompts a closer examination of how beginner-heavy teams build on prior work. Our atypical combination score-based results give one nuance: such teams draw from an unconventionally broader range of prior knowledge domains, as reflected in their use of unusual reference venue pairings. To further probe this, we examine two metrics: reference popularity (mean citations of a paper’s references; higher values reflect citing “hotter” work) and reference age (mean year gap between a paper and its references; lower values reflect more recent sources).
@@
Another key difference between beginners and later-career authors is productivity. Beginners publish fewer papers in their debut year—averaging $1.18$ in year 0 versus $1.68$ for career-age $1$ authors—a significant increase of +0.50+0.50 (Kruskal–Wallis test, $P<0.001$) that persists across subsequent years, with only modest annual increases afterward (maximum +0.13+0.13; Supplementary Figure S6). This structural difference suggests that beginners initially emphasize novelty over volume, echoing prior work showing that higher productivity can reduce disruption [56, 46, 45, 57, 58, 59]. Taken together, our results suggest that beginners tend to favor breadth over quantity in ways that are beneficial for disruption but may become challenging to sustain later in their careers.
Figure 3: Early-career collaborators are associated with higher disruption in beginner-heavy teams. We illustrate disruptions of pairwise combinations of beginner, early-career, and senior authors across different author ratios. The heatmap cells visualize median disruption percentiles, with darker shades indicating higher disruption. While teams heavier on beginners are progressively more disruptive, early-career collaborators are associated with better beginner disruption than senior collaborators. See Supplementary Tables S5–S7 for more details.
Figure 4: The most disruptive team compositions are beginner-heavy. We show the top 2020 most disruptive team compositions in our dataset (ranked by mean disruption percentiles; minimum $n\geq 10000$ per composition). Each bar represents a unique team composition, with emojis indicating the exact number of authors from each career stage and colors denoting their proportional representation in the teams. Beginner authors systematically dominate the most disruptive compositions.
Early-career and disruptive collaborators are linked to greater disruption in beginner-heavy teams
@@
Finally, we examine beginner ratios in the highly prevalent and highly disruptive team compositions in our dataset. We isolate high-frequency team compositions (each with $\geq 10,000$ papers and team size $\leq 40$) to identify $277$ unique compositions of beginner, early-career, and senior author counts. We rank these compositions by their mean disruption percentiles (Figure 4). Remarkably, we find that beginner-rich teams dominate: 6464% of the top $50$ compositions include equal or greater representation of beginners, and in 5252% of these, beginners outnumber all other career stages. These findings underscore the potential of beginner-heavy teams to drive scientific disruption.
Highly disruptive papers by beginner-heavy teams are highly cited
How well is disruptive work by beginner-heavy teams received by the academic community? Citations are commonly used as a proxy for academic impact [60]. Prior work documented a negative overall relationship between disruption scores and citation counts [61, 62], a result corroborated in our dataset. Unsurprisingly, we observe an overall negative association between the proportion of beginner authors in teams and their citation impact (Supplementary Figure S9).
However, this aggregate trend obscures a more nuanced citation dynamic among the most disruptive papers [61]. To uncover it, we first rank all papers by disruption scores and citation counts. Then, for each cumulative top-NN% of disruptive papers ($N=10$ to $100$, in steps of $10$), we compute the Spearman correlation between disruption and citation ranks within each beginner-ratio quartile (Q1–Q4). A positive correlation implies that high disruptiveness is rewarded with high citation impact, and vice versa.

### v2

@@ Abstract
Teams now drive most scientific advances, yet the impact of absolute beginners—authors with no prior publications—remains understudied. Analyzing over 29 million articles published between 1941 and 2020 across disciplines and team sizes, we uncover a near-universal and previously undocumented pattern: teams with a higher fraction of beginners are systematically more disruptive and innovative. Their contributions are linked to distinct knowledge-integration behaviors, including drawing on broader and less canonical prior work and producing more atypical recombinations. Collaboration structure further shapes outcomes: disruption is high when beginners work with early-career colleagues or with co-authors who have disruptive track records. Although disruption and citations are negatively correlated overall, highly disruptive papers from beginner-heavy teams are highly cited. These findings reveal a “beginner’s charm” in science, highlighting the underrecognized yet powerful value of beginner fractions in teams and suggesting actionable strategies for fostering a thriving ecosystem of innovation in science and technology.
@@
Disruption, Innovation, Career Age, Beginner Authors, Team Science
1 Introduction
@@
Imagine a research team with a high fraction of “beginner” authors—individuals with no prior publication. How likely is this team to disrupt its field? Intuitively, one might expect such teams to be unlikely engines of discovery. Scientific creativity rarely arises ex nihilo; theories of scientific and technological change emphasize discovery as an endogenous, cumulative process in which past knowledge enables future progress [fleming2001recombinant, schilling2011recombinant, fleming2010lone, baten2022novel, baten2021cues, baten2020creativity, kelty2025innovation, baten2024ai]. In this process, extended training and accumulated expertise allow researchers to stand, in Newton’s phrase, “on the shoulders of giants” [acemoglu2016innovation, mukherjee2017nearly, kuhn1997structure]. Indeed, combinatorial accounts of creativity emphasize atypical recombinations of such well-established knowledge components as the substrate of breakthroughs [wang2017bias, lee2015creativity]. Moreover, as fields advance, the “burden of knowledge” intensifies, requiring ever longer training to reach the frontier and leaving less time to extend it [jones2009burden, jones2011age]. By this logic, beginner authors—who have had little time to digest this burden—may be handicapped in their capacity for field-shifting innovation and disruption. If mastery is required before divergence, then beginner-heavy teams may lack the cognitive materials and experiential grounding to engineer novelty.
@@
Yet, there are equally compelling reasons to expect that beginner-heavy teams may, in fact, be well-positioned to disrupt scientific advancement. Youth is historically linked to novelty in science: they are less encumbered by entrenched paradigms, more open to new ideas, and more willing to take intellectual risks [rappa1993youth, packalen2019age]. Freed from institutionalized pressures—such as administrative burdens and funding constraints that often steer senior researchers toward conservative publication strategies [cohen1985revolution, spreng2021exploration, livan2019don, fortunato2018science, simonton2014wiley, higashide2024mid, spencer2017research, mcalpine2018identity, schimanski2018evaluation]—beginners may pursue bold ideas more readily. Moreover, deep disciplinary expertise can harden into attachment, making it harder for experienced scientists to selectively “unlearn” prevailing assumptions and embrace radically new ones [cui2022aging, merton1973sociology, candia2021quantifying, lin2025disruption]. From this perspective, beginners’ absence of enculturation may be a strength precisely because they have not yet fully assimilated the burden of knowledge. Indeed, many landmark discoveries originated as debut publications—such as Turing Awardee Yann LeCun’s early doctoral work, which laid the foundation for backpropagation [lecun85, lecun_publis_2025]. Extending this logic to the team level, beginner-heavy groups may be especially disruptive: their fresh cognitive repertoires, freedom from orthodoxy, and high risk tolerance may outweigh individual knowledge deficits, enabling breakthrough contributions more readily than in teams with fewer beginners.
@@
While recent studies have explored the aggregated effects of mean career age on team disruption [yang2024unveiling, zeng2021fresh], the role of the very youngest contributors—beginner authors—remains overlooked. In particular, it remains unclear how the disruptivity of teams changes as the fraction of beginners increases, and what underlying team dynamics and knowledge integration mechanisms drive these outcomes. To address this gap, we conduct a large-scale study of more than 2929M articles published between $1941$ and $2020$ across $19$ broad scientific disciplines. Our analysis draws on SciSciNet V2 [lin2023sciscinet, sciscinet-v2], one of the most comprehensive and widely used bibliometric data lakes. Using SciSciNet’s name-disambiguated author data [torvik2009author], we calculate career age as the number of years since an author’s first publication [frandsen2024defining, bradshaw2021fairer, cui2022aging], and classify authors into three career stages for our main results: beginner (0 years), early-career (1–10 years), and senior researchers (11 or more years). These thresholds reflect typical academic transitions, with the early stages encompassing graduate training and initial faculty appointments, and seniority being attained after a decade of publishing [vilela2023career, bonaccorsi2003age]. We also conduct robustness tests using a different choice of career stage thresholds: beginner (0 years), early-career (1–5 years), mid-career (6-10 years), and senior researchers (11 or more years) to confirm the generality of the results.
@@
We find strong evidence for a near-universal and previously undocumented pattern of “beginner’s charm”: an increased presence of beginner authors in teams is associated with higher scientific disruption and innovation. This association holds across disciplines, time periods, and team sizes, suggesting it reflects a general feature of how science advances rather than a peculiarity of specific fields or historical moments.
@@ 2 Results
2.1 Teams with higher beginner-author ratios are more disruptive and innovative
@@
As our primary team performance metric, we use the disruption score, originally developed to distinguish between consolidating and destabilizing contributions in science and technology [funk2017dynamic]. This score ranges from 1-1 to +1+1, capturing papers that reinforce or shift scientific trajectories, respectively (see Materials and Methods). Since its introduction, the disruption score has been widely adopted and validated across domains [wu2019large, lin2023sciscinet, leibel2024we, chu2021slowed, park2023papers, wei2023quantifying], providing a robust lens for answering our research queries.
@@
We find that papers with a higher proportion of beginner authors tend to be more disruptive: in the overall dataset, the beginner ratio is positively correlated with a paper’s disruption percentile (Pearson’s $r=0.11$, Holm-adjusted $P<10^{-6}$), while early-career ratio shows a small positive association ($r=0.002$, $P<10^{-6}$), and senior-author ratio shows a negative association (r=0.08r=-0.08, $P<10^{-6}$).
@@
While the overall correlations are informative, they obscure the role of team size, since beginner ratios are mechanically constrained by the number of authors and disruption itself is known to vary with team size [wu2019large]. To address this, we stratify the analysis by team size and recompute the associations (Figure 1). The results are strikingly monotonic across team sizes: beginner ratios show uniformly strong positive correlations with disruption (Kendall’s $\tau$ up to 1.0 with narrow 95% CIs, $P<0.001$ in larger teams), senior ratios show consistently negative associations (often $\tau=-1.0$, $P<0.001$), and early-career ratios display weaker but positive effects that reach significance in larger teams (Supplementary Table S1; all $P$-values are Benjamini–Hochberg FDR-corrected).
@@
Prior studies found that larger teams tend to produce less disruptive work [wu2019large], and we replicate this pattern in teams with low beginner ratios. Interestingly, however, the positive effect of beginner ratio on disruption is especially pronounced in larger teams, where the usual size penalty attenuates (Supplementary Figure S1). As a result, beginner-heavy large teams approach the disruption levels of beginner-heavy small teams. One possible explanation is that a higher beginner ratio may counteract the convergence pressures of large teams—such as reduced risk-taking [christensen2011innovators], resistance to external perspectives [minson2012cost], and neutralized viewpoints [greenstein2016open]—enabling a shift from consolidating “safe bets” to more exploratory, high-variance work [paulus2013understanding, lakhani2013prize].
The positive association between beginner share and disruption is robust across all eight decades and in all 19 broad academic disciplines in the corpus ($r>0$, $P<0.001$ throughout; Supplementary Tables S2-S3). Decade-specific analyses show the positive association to hold true across team sizes ($r>0$, $P<10^{-6}$ throughout; Supplementary Figure S2). For discipline-specific analysis, we leverage SciSciNet’s field-of-study annotations derived from the Microsoft Academic Graph (MAG), which organizes research into a hierarchical taxonomy of fields [lin2023sciscinet, sinha2015overview]. In particular, we use SciSciNet’s 19 Level-0 (top-level) field annotations as broad disciplines, allowing each paper to be associated with one or more disciplines. We include all papers associated with a given discipline, such that each multidisciplinary paper contributes to multiple disciplines. Across all 19 disciplines, the positive association between beginner share and disruption remains robust across team sizes (Supplementary Figure S3). Collectively, these patterns indicate that disruption is robustly elevated in teams with a higher fraction of beginners.
To verify that our findings are not artifacts of the disruption score metric, we employ a secondary metric of atypical combination scores (see Materials and Methods) [jones2009burden, uzzi2013atypical, guimera2005team, jones2008multi]. This score conceptualizes discovery as innovative knowledge-recombination: for each paper, all pairs of reference venues are scored for atypicality against a degree- and time-preserving null, and the median $z$-score of all these pairs is taken as the paper’s atypical combination score. A score below 0 represents innovative combinations of existing knowledge, and vice versa [uzzi2013atypical]. Unlike disruption (a downstream impact measure), the atypical combination score captures innovation from a paper’s inputs.
Figure 2 shows the cumulative distributions of atypical combination scores for different beginner author-ratio quartiles (Q1: 0–0.25, Q2: 0.25–0.50, Q3: 0.50–0.75, Q4: 0.75–1.00). As the share of beginner authors increases, the cumulative distributions shift toward lower values, indicating more innovative combinations. The ECDFs of the atypical combination scores differ significantly between every quartile pair (two-sample Kolmogorov–Smirnov test; Holm-adjusted $P<10^{-300}$ for all comparisons; largest contrast Q1 vs Q4: $D=0.142$). Supplementary Figure S4 shows that higher proportions of beginner authors are associated with lower atypicality score percentiles (higher innovation). In contrast, higher senior-author ratios are associated with higher atypicality score percentiles (lower innovation). Together, these findings corroborate and validate our disruption-based results that beginner-heavy teams are more innovative than beginner-light ones.
@@
The marked difference in disruption and innovation between beginner-heavy and other teams prompts a closer examination of how beginner-heavy teams build on prior work. Our atypical combination score-based results reveal one nuance: such teams draw from an unconventionally broader range of prior knowledge domains, as reflected in their use of unusual reference-venue pairings. To further probe this, we examine two metrics: reference popularity (mean citations of a paper’s references; higher values reflect citing “hotter” work) and reference age (mean year gap between a paper and its references; lower values reflect more recent sources).
@@
Another key difference between beginners and later-career authors is productivity. Beginners publish fewer papers in their debut year—averaging $1.18$ in year 0 versus $1.68$ for career-age $1$ authors—a significant increase of +0.50+0.50 (Kruskal–Wallis test, $P<0.001$) that persists across subsequent years, with only modest annual increases afterward (maximum +0.13+0.13; Supplementary Figure S6). This structural difference suggests that beginners initially emphasize novelty over volume, echoing prior work showing that higher productivity can reduce disruption [li2024productive, park2023papers, chu2021slowed, li2022bibliographic, michalska2017and, li2025quantifying]. Taken together, our results suggest that beginners tend to favor breadth over quantity in ways that are beneficial for disruption but may become challenging to sustain later in their careers.
Figure 3: Early-career collaborators are associated with higher disruption in beginner-heavy teams. We illustrate disruptions of pairwise combinations of beginner, early-career, and senior authors across different author ratios. The heatmap cells visualize median disruption percentiles, with darker shades indicating higher disruption. While teams heavier on beginners are progressively more disruptive, early-career collaborators are associated with better beginner disruption than senior collaborators. See Supplementary Tables S5-S7 for more details.
Figure 4: The most disruptive team compositions are beginner-heavy. We show the top $10$ most disruptive team compositions in our dataset (ranked by mean disruption percentiles; minimum $n\geq 10,000$ per composition). Each bar represents a unique team composition, with emojis indicating the exact number of authors from each career stage and colors denoting their proportional representation in the teams. Beginner authors systematically dominate the most disruptive compositions.
2.2 Early-career and disruptive collaborators are linked to greater disruption in beginner-heavy teams
@@
Finally, we examine beginner ratios in the highly prevalent and highly disruptive team compositions in our dataset. We isolate high-frequency team compositions (each with $\geq 10,000$ papers and team size $\leq 40$) to identify $103$ unique compositions of beginner, early-career, and senior author counts. We rank these compositions by their mean disruption percentiles (Figure 4). Remarkably, beginner-rich teams remain prominent: among the top $50$ most disruptive compositions, $24$ have at least as many beginners as other career stages, and in $15$ cases, beginners outnumber all other career stages combined. Crucially, this effect is not a fluke: each composition reflects at least $10,000$ papers over eight decades, revealing a robust, recurring beginner advantage with consistently high disruption.
2.3 Highly disruptive papers by beginner-heavy teams are highly cited
How well is disruptive work by beginner-heavy teams received by the academic community? Citations are commonly used as a proxy for academic impact [sinatra2016quantifying]. Prior work documented a negative overall relationship between disruption scores and citation counts [li2024breaking, zeng2023disruptive], a result corroborated in our dataset. Unsurprisingly, we observe an overall negative association between the proportion of beginner authors in teams and their citation impact (Supplementary Figure S9).
However, this aggregate trend obscures a more nuanced citation dynamic among the most disruptive papers [li2024breaking]. To uncover it, we first rank all papers by disruption scores and citation counts. Then, for each cumulative top-NN% of disruptive papers ($N=10$ to $100$, in steps of $10$), we compute the Spearman correlation between disruption and citation ranks within each beginner-ratio quartile (Q1–Q4). A positive correlation implies that high disruptiveness is rewarded with high citation impact, and vice versa.